suppose that the average outstanding credit card balance for young professionals is Php 11,200 with standard deviation of Php 2,600. In simple random sample of 150 young professionals, what is the probability that the mean outstanding credit card balance does not exceeds Php 12, 300?

1
Expert's answer
2022-03-14T14:46:07-0400

"\mu=11200,\sigma=2600,n=150"

"P(x\u226412300)??" "Z=(x-μ)/(σ⁄\sqrt{n})"

Therefore "P(Z≤(12300-11200)/(2600⁄\sqrt{150})"

"=P(Z\u22645.18)=0.9999"